Find acute angle between intersecting curves. Write normal line, tangent line for each curves at the intersec tion point.


y = x ^ 3 , x = y ^ 2


1
Expert's answer
2021-10-18T15:13:41-0400

Solution:

"y=x^3\ ...(i)
\\x=y^2\ ...(ii)"

Put (ii) in (i)

"y=(y^2)^3
\\ \Rightarrow y=y^6
\\ \Rightarrow y-y^6=0
\\ \Rightarrow y(1-y^5)=0
\\ \Rightarrow y=0, y=1"

When "y=0,x=0"

When "y=1,x=1"

So, the points of intersection are (0,0), (1,1).

Now, for (0,0):

"y=x^3
\\ \Rightarrow y'=3x^2
\\ \Rightarrow y'(0,0)=3(0)=0=m_1"

"x=y^2
\\\Rightarrow y^2=x
\\\Rightarrow 2y.y'=1
\\\Rightarrow y'=\dfrac1{2y}
\\\Rightarrow y'(0,0)=\dfrac10=\infty=m_2"

Now, "\tan \theta=|\dfrac{m_1-m_2}{1+m_1m_2}|=|\dfrac{0-\infty}{1+0.\infty}|=\infty=not\ defined"

Thus, rejecting this case.

Next, for (1,1):

"y=x^3
\\ \Rightarrow y'=3x^2
\\ \Rightarrow y'(1,1)=3(1)^2=3=m_1"

"x=y^2
\\\Rightarrow y^2=x
\\\Rightarrow 2y.y'=1
\\\Rightarrow y'=\dfrac1{2y}
\\\Rightarrow y'(1,1)=\dfrac1{2(1)}=\dfrac12=m_2"

Now, "\tan \theta=|\dfrac{m_1-m_2}{1+m_1m_2}|=|\dfrac{3-\dfrac12}{1+3\times\dfrac12}|=1"

So, "\theta=45\degree"